The power drawn by a resistive coil made of copper wire is 220 W at 100 V and 20 °C. Calculate the power consume by the coil at 115 V and 100°C. The temperature coefficient of resistance at 20°C is 0.00393. (include step-by-step procedures)


1
Expert's answer
2021-10-04T02:29:06-0400

1) Find resistance at 20°C:


"R=\frac{V^2}{P}=45.5\Omega."

2) Find resistance at 100°C:


"R_1=R(1+\alpha \Delta T)=59.7\Omega."

3) Find the power consumed at 115 V:


"P_1=\frac{V_1^2}{R_1}=222\text W."